# Transition to inverse cascade in turbulent rotating convection in absence of the large-scale vortex

### v1

@@ Abstract
Turbulent convection under strong rotation can develop an inverse cascade of kinetic energy from smaller to larger scales. In the absence of an effective dissipation mechanism at the large scales, this leads to the pile-up of kinetic energy at the largest available scale, yielding a system-wide large-scale vortex (LSV). Earlier works have shown that the transition into this state is abrupt and discontinuous. Here, we study the transition to the inverse cascade in the case where the inverse energy flux is dissipated before it reaches the system scale, suppressing the LSV formation. We demonstrate how this can be achieved in direct numerical simulations by using an adapted form of hypoviscosity on the horizontal manifold. We find that in the absence of the LSV, the transition to the inverse cascade becomes continuous. This shows that it is the interaction between the LSV and the background turbulence that is responsible for the observed discontinuity. We furthermore show that the inverse cascade in absence of the LSV has a more local signature compared to the case with LSV.
@@ 1 Introduction
Many of the convective flows that are encountered in the geo- and astrophysical settings are also subject to strong rotation, typically due to rotation of the celestial body itself, which can dramatically alter the flow (Ecke & Shishkina, 2023). When rotation is strong, it suppresses velocity fluctuations along the rotation axis, rendering the flow quasi-two-dimensional (quasi-2D). As a result of the quasi-two-dimensionalisation, the flow can develop an inverse cascade of kinetic energy, where part of the kinetic energy is transferred from smaller length scales to larger scales, opposite from the conventional direct cascade in regular 3D turbulence. In absence of an effective dissipation mechanism at the large scales, this energy will pile up at the largest available scale. This condensate of kinetic energy at the largest scales manifests itself in the form of a strong system-wide vortex, which in rotating convection has been coined the large-scale vortex or LSV (Favier et al., 2014; Guervilly et al., 2014; Julien et al., 2012; Rubio et al., 2014). Recently, it has been shown that the transition from the forward cascading 3D-like state to the inverse cascading quasi-2D LSV state in rotating convection is, remarkably, an abrupt and discontinuous transition (Favier et al., 2019; de Wit et al., 2022a). This is in close similarity with the condensate transition observed in other quasi-2D flow systems, such as thin-layer turbulence (van Kan & Alexakis, 2019; de Wit et al., 2022b) and forced rotating turbulence (Yokoyama & Takaoka, 2017), see also Alexakis & Biferale (2018), Alexakis (2023) and van Kan (2024) for recent reviews.
@@
In some cases, however, the inverse cascade is dissipated before it reaches the largest system scale. In that case, no large-scale condensate forms. This is common in systems where there is a very large separation between the typical scales at which the energy is injected and the total system scale, such that any ostensibly small friction mechanism that is active at the large scales (e.g. some boundary friction) will deplete the inverse cascade before it reaches the system scale. In direct numerical simulations (DNS) this is typically achieved by applying a hypoviscosity, which is a dissipation mechanism that is artificially restricted to the largest scales, to mimic a large degree of scale separation. Earlier works in other quasi-2D turbulent flow systems have revealed that the absence of the condensate can drastically alter the inverse cascade transition (Seshasayanan et al., 2014; Benavides & Alexakis, 2017; Pestana & Hickel, 2019; van Kan & Alexakis, 2020). In this work, we study the inverse cascade transition in the natural flow system of rotating convection under the influence of a large-scale friction that prevents the formation of the LSV. This allows us to disentangle the influence of the LSV from the dynamics of the underlying turbulence in regards to the inverse cascade transition.
@@ 2 Numerical methods
$\textrm{Ra}\equiv\frac{g\alpha\Delta TH^{3}}{\nu\kappa},\qquad\textrm{Ek}\equiv\frac{\nu}{2\Omega H^{2}},\qquad\textrm{Pr}\equiv\frac{\nu}{\kappa}.$ (1)
Here, $g$ denotes the gravitational acceleration, $H$ is the domain height, while the fluid properties $\alpha$, $\nu$ and $\kappa$ represent the thermal expansion coefficient, kinematic viscosity and thermal diffusivity of the fluid, respectively.
#runs Ra Pr Ek $\nu_{\alpha}$ $\alpha$ $L/H$ $N_{x}\times N_{y}\times N_{z}$ With hypoviscosity 22 $[2\times 10^{6}-5\times 10^{7}]$ 1 $10^{-4}$ $1.5\times 10^{2}$ 4 2.235 $257\times 257\times 129$ With hypoviscosity (wide) 22 $[2\times 10^{6}-5\times 10^{7}]$ 1 $10^{-4}$ $3.8\times 10^{-2}$ 1 4.470 $513\times 513\times 129$ Without hypoviscosity 46 $[2\times 10^{6}-5\times 10^{7}]$ 1 $10^{-4}$ 2.235 $257\times 257\times 129$ covered in de Wit et al. (2022a)
@@
$\frac{\partial\bm{u}}{\partial t}+\left(\bm{u}\cdot\bm{\nabla}\right)\bm{u}+\frac{\left(\textrm{Pr}/\textrm{Ra}\right)^{1/2}}{\textrm{Ek}}\bm{e}_{z}\times\bm{u}$ $=-\bm{\nabla}p+T\bm{e}_{z}+\left(\frac{\textrm{Pr}}{\textrm{Ra}}\right)^{1/2}\Delta\bm{u}-\underbrace{\nu_{\alpha}\Delta_{h}^{-\alpha}\bm{u}_{h}}_{\text{hypoviscosity}},$ (2a)
$\frac{\partial T}{\partial t}+\left(\bm{u}\cdot\bm{\nabla}\right)T$ $=\frac{1}{\left(\textrm{Ra}\textrm{Pr}\right)^{1/2}}\nabla^{2}T,$ (2b)
@@
Notice that in the last term of equation (2a) we have included the hypoviscosity. This is the large-scale friction term that takes care of dissipating the inverse energy flux. However, in a doubly periodic domain like in the (rotating) convection system at hand, we need to be somewhat careful how we can define an effective hypoviscosity, as we cannot simply use the isotropic inverse Laplacian that is commonly applied in triply periodic domains. Instead, motivated by the fact that we want to dissipate the inverse cascade in the 2D horizontal manifold, we apply a 2D variant of the hypoviscosity that acts on the horizontal velocity and the horizontal inverse Laplacian, applied across the full height of the domain. This is conveniently computed from a 2D Fourier transform in the horizontal periodic directions $\mathcal{F}_{h}\{...\}$, taking the horizontal velocity $\bm{u}_{h}\equiv u_{x}\bm{e}_{x}+u_{y}\bm{e}_{y}$ from the real space $\bm{u}_{h}(x,y,z,t)$ to the Fourier space $\hat{\bm{u}}_{h}(k_{x},k_{y},z,t)$. Then, we can compute the hypoviscous term as
@@
We consider a fluid with $\textrm{Pr}=1$, rotating at $\textrm{Ek}=10^{-4}$ and vary Ra over the range in which the flow is known to transition into and out of the inverse cascade regime, in agreement with the parameter range considered in de Wit et al. (2022a) for the case with LSV (without hypoviscosity). See table 1 for the full set of input parameters considered in this work. The power $\alpha$ and strength $\nu_{\alpha}$ of the hypoviscosity are tuned in such a way that the full inverse energy flux is dissipated before it reaches the system scale, while simultaneously ensuring that its action remains mostly restricted to the largest scales. We consider two series of runs with different horizontal domain sizes. The aspect ratio of the first series is chosen in agreement with de Wit et al. (2022a), while the second series of runs doubles the horizontal extent in order to increase the scale separation between the system scale and the scales that dominate the energy injection. The hypoviscosity parameters are kept constant throughout each series of runs. The grid is uniform in the horizontal directions, but non-uniform in the vertical direction, narrowing the grid points close to the top and bottom to accurately resolve the boundary layers. The resolution $N_{x}\times N_{y}\times N_{z}$ is chosen with respect to the Kolmogorov scale $\eta$, ensuring that the grid spacing never exceeds $2\eta$, which is well below the threshold of $4\eta$ that is obtained for this code in Verzicco & Camussi (2003).
@@ 3 Rotating convection with large-scale friction
Figure 1: (a) Instantaneous kinetic energy spectra $E(k)$ for the cases with $\textrm{Ra}=8\times 10^{6}$ with and without hypoviscosity and for different horizontal domain sizes (see table 1). (b,c) The corresponding snapshots of kinetic energy $\tfrac{1}{2}|\bm{u}|^{2}$ at the mid-plane $z=H/2$.
The hypoviscosity is applied in order to introduce a friction mechanism at the large scales that prevents the formation of the LSV. To show the influence of the hypoviscosity on the flow we consider the kinetic energy spectrum
@@
    The hypoviscosity needs to be able to dissipate the full inverse energy flux before it condensates into an LSV. In practice, this means that the hypoviscosity must be strong enough such that the spectral energy contained in the smallest (system scale) wavenumber remains smaller than the energy contained in the next wavenumbers, that is $E(k=2\pi/L)<E(k>2\pi/L)$.
@@
    The action of the hypoviscosity needs to remain restricted to the smallest wavenumbers. We ensure this by checking that the spectral slope at small wavenumbers does not exceed $\mathrm{d}\log E(k)/\mathrm{d}\log k<2\alpha$, such that the action of the hypoviscosity is indeed dominated by the smallest wavenumbers.
Then, if there is sufficient scale separation between the large scales where the hypoviscosity acts and the scales that dominate the energy injection, conservation of energy ensures that the rate of energy dissipation by the hypoviscosity will match exactly with the inverse energy flux, independent of the precise values of the hypoviscosity parameters, as long as the aforementioned conditions are met. This is verified numerically for an exemplary case in appendix A.
Then, considering fluctuation fields that are dominated by intermediate- and small-scale structures, they remain unaffected by the presence of hypoviscosity, see for example figure 2.
Figure 2: Instantaneous snapshots of the temperature fluctuation field $T-\tilde{T}(z)$, where $\tilde{T}(z)$ is the average temperature profile, for the cases with $\textrm{Ra}=8\times 10^{6}$ without hypoviscosity (a) and with hypoviscosity (b).
@@ 4 Transition towards inverse cascade
With the hypoviscosity in place and tuned, we are well-equipped to study the transition towards the inverse energy cascade under the influence of large-scale friction, in absence of the LSV. To that extent, we consider the strength of the inverse energy flux $\epsilon_{\textrm{inv}}$ as we vary the control parameter Ra throughout the regime where the inverse cascade is known to occur.
As we argued, when the action of the hypoviscosity is correctly tuned, the energy dissipated by the hypoviscosity coincides with the inverse energy flux. We can thus measure the mean inverse flux in this case as
$\epsilon_{\textrm{inv}}^{\textrm{(hypo)}}\equiv\left\langle\nu_{\alpha}\bm{u}_{h}\cdot\Delta_{h}^{-\alpha}\bm{u}_{h}\right\rangle,$ (6)
where we use delimited-⟨$\langle$..$\rangle$ … ⟩ to denote the average over the full spatial domain and time.
@@
$\epsilon_{\textrm{inv}}^{\textrm{(no-hypo)}}\equiv\sum_{Q}T_{\textrm{3D}}(K=\tfrac{2\pi}{L},Q)+T_{\textrm{2D}}(K=\tfrac{2\pi}{L},Q)$ (7)
@@
where the overbar denotes vertical averaging ¯$\mkern 1.5mu\overline{\mkern-1.5mu\vphantom{a}...\mkern-1.5mu}\mkern 1.5mu\equiv\tfrac{1}{H}\int_{0}^{H}...\mathrm{d}z$ and with $\bm{u}^{\textrm{2D}}\equiv\mkern 1.5mu\overline{\mkern-1.5mu\bm{u}_{h}\mkern-1.5mu}\mkern 1.5mu$ and $\bm{u}^{\textrm{3D}}\equiv\bm{u}-\bm{u}^{\textrm{2D}}$ and subscript $K$ denoting Fourier-filtering at mode $K$.
@@
Figure 3: The total inverse flux of kinetic energy $\epsilon_{\textrm{inv}}$ normalized by the total energy flux $\epsilon=(\textrm{Nu}-1)/\sqrt{\textrm{Pr}\textrm{Ra}}$ as a function of Ra for $\textrm{Ek}=10^{-4}$ and $\textrm{Pr}=1$ for the different series of runs with and without hypoviscosity. Panel (b) is a zoom of panel (a) that focuses on the cases with hypoviscosity. It shows that while the transition to the inverse cascade without hypoviscosity is discontinuous and hysteretic (lower hysteretic branch is depicted by the dotted line), the transition for the case with hypoviscosity is continuous.
@@
The results are shown in figure 3. While the transition to the inverse cascade regime without large-scale friction, in presence of the LSV, is known to be discontinuous and hysteretic (Favier et al., 2019; de Wit et al., 2022a), we find that this behaviour is entirely different for the case with large-scale friction, in absence of the LSV. Here, we observe the transition in this case is in fact smooth and the discontinuity disappears. This is in agreement with what is found in other hypoviscous quasi-2D systems (Seshasayanan et al., 2014; Benavides & Alexakis, 2017; Pestana & Hickel, 2019; van Kan & Alexakis, 2020). We furthermore find that the strength of the inverse cascade in absence of the LSV is much smaller than for the cases that have an LSV.
@@
A point of attention is raised by the observation that the inverse cascade does not seem to decay to zero for large Ra for the case with hypoviscosity. However, we argue that this is an effect of insufficient scale separation between the large scales at which the hypoviscosity acts and the scales that dominate the energy injection, such that the hypodissipation not only measures the inverse energy flux, but also the fraction of energy that is directly injected at these large scales. The energy injection is known to be dominated by features of typical size $\mathcal{L}_{c}\equiv 4.8\textrm{Ek}^{1/3}H$, the most unstable wavelength at onset of convection (Chandrasekhar, 1961), but the self-organizing nature of the forcing injects energy at a broad range of scales around this dominant scale. The runs without hypoviscosity (red line) were performed at $L\approx 10\mathcal{L}_{c}$ and the runs with hypoviscosity (blue line) employ the same aspect ratio in order to have a one-to-one comparison. However, in order to increase the scale separation, we have added the series of runs in the wider domain $L\approx 20\mathcal{L}_{c}$, which also allows for a less aggressive exponent of the hypoviscosity (see table 1). Indeed, we notice that the measured energy flux that is dissipated by the hypoviscosity remains mostly unchanged, but reaches a smaller value at large Ra for the wider domain. We thus expect that the inverse flux will decay to values close to zero in the limit of increasingly large domain size as the scales become more cleanly separated.
@@
Figure 4: Time-averaged kinetic energy transport maps from 3D (a,c) and 2D (b,d) modes $Q$ to 2D modes $K$, i.e. $T_{3D}(K,Q)$ and $T_{2D}(K,Q)$, respectively, for the cases with $\textrm{Ra}=8\times 10^{6}$ without hypoviscosity (a,b) and with hypoviscosity (c,d).
@@
To further understand the influence of the presence/absence of the LSV on the nature of the inverse cascade, we also consider the scale-to-scale kinetic energy transfer maps. By looking at $T_{\textrm{3D}}(K,Q)$ and $T_{\textrm{2D}}(K,Q)$, we can study how (baroclinic and barotropic) kinetic energy is transferred from donating wavenumber $Q$ to receiving wavenumber $K$. The results are provided in figure 4 for an exemplary case with and without hypoviscosity. It shows that in presence of the LSV (upper panels), the upscale energy transfer is highly non-local, with a direct transfer from all wavenumbers $Q$ into the largest wavenumber $K=1$, as is also known from Rubio et al. (2014); Aguirre Guzmán et al. (2020); de Wit et al. (2022a). In absence of the LSV (lower panels), on the contrary, we find that the upscale energy transfer has a much more local signature, with the energy transfer concentrating more around the diagonal of the map, indicating transfer between more like-sized scales.
Finally, we can also study what the influence of the presence/absence of the LSV is on the total thermal throughput of the convective system, quantified through the Nussel number Nu, given by

### v2

@@ Abstract
Turbulent convection under strong rotation can develop an inverse cascade of kinetic energy from smaller to larger scales. In the absence of an effective dissipation mechanism at the large scales, this leads to the pile-up of kinetic energy at the largest available scale, yielding a system-wide large-scale vortex (LSV). Earlier works have shown that the transition into this state is abrupt and discontinuous. Here, we study the transition to the inverse cascade at Ekman number $\textrm{Ek}=10^{-4}$ and using stress-free boundary conditions, in the case where the inverse energy flux is dissipated before it reaches the system scale, suppressing the LSV formation. We demonstrate how this can be achieved in direct numerical simulations by using an adapted form of hypoviscosity on the horizontal manifold. We find that in the absence of the LSV, the transition to the inverse cascade becomes continuous. This shows that it is the interaction between the LSV and the background turbulence that is responsible for the earlier observed discontinuity. We furthermore show that the inverse cascade in absence of the LSV has a more local signature compared to the case with LSV.
@@ 1 Introduction
Many of the convective flows that are encountered in the geo- and astrophysical settings are also subject to strong rotation, typically due to rotation of the celestial body itself, which can dramatically alter the flow (Ecke & Shishkina, 2023). When rotation is strong, it suppresses variations of the velocity along the rotation axis, promoting quasi-two-dimensional (quasi-2D) flow organization. As a result of the quasi-two-dimensionalisation, the flow can develop a partial inverse cascade of kinetic energy, where a fraction of the kinetic energy is transferred from smaller length scales to larger scales, opposite from the direct cascade in regular 3D turbulence. In absence of an effective dissipation mechanism at the large scales, this energy will pile up at the largest available scale. This condensate of kinetic energy at the largest scales manifests itself in the form of a strong system-wide vortex, which in rotating convection has been coined the large-scale vortex or LSV (Favier et al., 2014; Guervilly et al., 2014; Julien et al., 2012; Rubio et al., 2014). Recently, it has been shown that the transition from the forward cascading 3D-like state to the inverse cascading quasi-2D LSV state in rotating convection is, remarkably, an abrupt and discontinuous transition (Favier et al., 2019; de Wit et al., 2022a). In particular, the ‘high-thermal-forcing’ transition, where the rotational constraint of the flow breaks down to due finite rotation effects, was shown to be strongly hysteretic. This is in close similarity with the condensate transition observed in other quasi-2D flow systems, such as thin-layer turbulence (van Kan & Alexakis, 2019; de Wit et al., 2022b) and forced rotating turbulence (Yokoyama & Takaoka, 2017), see also Alexakis & Biferale (2018), Alexakis (2023) and van Kan (2024) for recent reviews.
@@
In some cases, however, the inverse cascade can be affected by a dissipation mechanism at large scales. If such a mechanism is sufficiently strong, it can dissipate the inverse energy flux entirely before it reaches the largest system scale and thereby suppress the formation of the large-scale condensate. A prime example is boundary friction, which has been shown to be able to partially or fully suppress the formation of the LSV in rotating convection (Kunnen et al., 2016; Julien et al., 2016; Plumley et al., 2016; Aguirre Guzmán et al., 2020). As system size increases, the larger scale separation between injection scales and the system scales allows a larger range for such a dissipation mechanism to be active, so that even an ostensibly small dissipation mechanism can eventually dissipate the inverse flux entirely before it reaches the system scale. In direct numerical simulations (DNS) this scenario is typically studied by applying a hypoviscosity: a dissipation mechanism that is artificially restricted to the largest scales, in order to mimic a large degree of scale separation. Earlier works in other quasi-2D turbulent flow systems have revealed that such large-scale friction can drastically alter the inverse cascade transition (Seshasayanan et al., 2014; Benavides & Alexakis, 2017; Pestana & Hickel, 2019; van Kan & Alexakis, 2020), where in particular, the discontinuity in the transition can be partially or entirely suppressed when the large-scale condensate is absent. In this work, we study the inverse cascade transition in the natural flow system of rotating convection under the influence of a large-scale friction that prevents the formation of the LSV. This allows us to disentangle the effect of the LSV from the dynamics of the underlying turbulence to understand its influence on the inverse cascade transition.
@@ 2 Numerical methods
$\textrm{Ra}\equiv\frac{g\gamma\Delta TH^{3}}{\nu\kappa},\qquad\textrm{Ek}\equiv\frac{\nu}{2\Omega H^{2}},\qquad\textrm{Pr}\equiv\frac{\nu}{\kappa}.$ (1)
Here, $g$ denotes the gravitational acceleration, $H$ is the domain height, while the fluid properties $\gamma$, $\nu$ and $\kappa$ represent the thermal expansion coefficient, kinematic viscosity and thermal diffusivity of the fluid, respectively.
@@
$\frac{\partial\bm{u}}{\partial t}+\left(\bm{u}\cdot\bm{\nabla}\right)\bm{u}+\frac{\left(\textrm{Pr}/\textrm{Ra}\right)^{1/2}}{\textrm{Ek}}\bm{e}_{z}\times\bm{u}$ $=-\bm{\nabla}p+T\bm{e}_{z}+\left(\frac{\textrm{Pr}}{\textrm{Ra}}\right)^{1/2}\Delta\bm{u}-\underbrace{\nu_{\alpha}\Delta_{h}^{-\alpha}\bm{u}_{h}}_{\begin{subarray}{c}\text{adapted}\\ \text{hypoviscosity}\end{subarray}},$ (2a)
$\frac{\partial T}{\partial t}+\left(\bm{u}\cdot\bm{\nabla}\right)T$ $=\frac{1}{\left(\textrm{Ra}\textrm{Pr}\right)^{1/2}}\Delta T,$ (2b)
@@
Notice that in the last term of equation (2a) we have included the adapted hypoviscosity. This is the large-scale friction term that takes care of dissipating the inverse energy flux. In a doubly periodic domain like in the (rotating) convection system at hand, we need to be somewhat careful how we can define an effective hypoviscosity, as we cannot simply use the isotropic inverse Laplacian that is commonly applied in triply periodic domains. Instead, motivated by the aim of dissipating the inverse cascade that resides in the 2D horizontal manifold, we apply a 2D variant of the hypoviscosity that acts on the horizontal velocity and the horizontal inverse Laplacian, applied across the full height of the domain. This is conveniently computed from a 2D Fourier transform in the horizontal periodic directions $\mathcal{F}_{h}\{$...\}, taking the horizontal velocity $\bm{u}_{h}\equiv u_{x}\bm{e}_{x}+u_{y}\bm{e}_{y}$ from the real space $\bm{u}_{h}(x,y,z,t)$ to the Fourier space $\hat{\bm{u}}_{h}(k_{x},k_{y},z,t)$. Then, we can compute the hypoviscous term as
@@
A different approach for suppression of the LSV in rotating convection is put forward by Oliver et al. (2023), based on a modification of the non-linear term. In appendix 6, we compare the different approaches for suppression of the LSV for an exemplary case.
We consider a fluid with $\textrm{Pr}=1$, rotating at $\textrm{Ek}=10^{-4}$ and vary Ra over the range in which the flow is known to transition into and out of the inverse cascade regime, in agreement with the parameter range considered in de Wit et al. (2022a) for the case with LSV (without hypoviscosity). See table 1 for the full set of input parameters considered in this work. The power $\alpha$ and strength $\nu_{\alpha}$ of the hypoviscosity are tuned in such a way that the full inverse energy flux is dissipated before it reaches the system scale, while simultaneously ensuring that its action remains mostly restricted to the largest scales (see also next section). We consider three series of runs with different horizontal domain sizes. The aspect ratio of the first series is chosen in agreement with de Wit et al. (2022a), employing $L=10\mathcal{L}_{c}$, with $\mathcal{L}_{c}\equiv 4.8\textrm{Ek}^{1/3}H$, the most unstable wavelength at onset of convection (Chandrasekhar, 1961). The second and third series of runs progressively double the horizontal extent in order to increase the scale separation between the system scale and the scales that dominate the energy injection. The hypoviscosity parameters are kept constant throughout each series of runs. The grid is uniform in the horizontal directions, but non-uniform in the vertical direction, narrowing the grid close to the top and bottom to accurately resolve the boundary layers. The resolution $N_{x}\times N_{y}\times N_{z}$ is chosen with respect to the Kolmogorov scale $\eta$, ensuring that the grid spacing never exceeds $2\eta$, which is well below the threshold of $4\eta$ that is obtained for this code in Verzicco & Camussi (2003).
#runs Ra Pr Ek $\nu_{\alpha}$ $\alpha$ $L/H$ $N_{x}\times N_{y}\times N_{z}$ With hypoviscosity I 22 $2\times 10^{6}-5\times 10^{7}]$ 1 $10^{-4}$ $1.5\times 10^{2}$ 4 2.235 $257\times 257\times 129$ With hypoviscosity II 22 $2\times 10^{6}-5\times 10^{7}]$ 1 $10^{-4}$ $3.8\times 10^{-2}$ 1 4.470 $513\times 513\times 129$ With hypoviscosity III 22 $2\times 10^{6}-5\times 10^{7}]$ 1 $0^{-4} 1.1$$1\times 10^{-2}$ 0.75 8.940 $1025\times 1025\times 129$ Without hypoviscosity 46 $2\times 10^{6}-5\times 10^{7}]$ 1 $10^{-4}$ N/A N/A 2.235 $257\times 257\times 129$ covered in de Wit et al. (2022a)
Simulations are initialized with zero velocity and a linear temperature profile connecting the top and bottom boundaries, superposed by a tiny statistical perturbation to initiate convection. Statistics are then collected after 1000 convective time units when a statistically steady state is reached at all scales.
@@ 3 Rotating convection with large-scale friction
Analogous to many other quasi-2D flow systems, rotating convection can give rise to a split cascade scenario (Alexakis & Biferale, 2018). In this scenario, from the intermediate scales at which energy is injected, part of the energy is cascaded to small scales as in regular 3D turbulence, while another fraction of the energy is cascaded from the intermediate scales to larger scales, owing to the quasi-2-dimensionalization. This gives rise to a net forward flux of energy at smaller scales, but a net inverse flux of energy at larger scales. While the forward flux is balanced by regular viscous dissipation, we shall here balance the inverse flux of energy with hypoviscosity at the large scales, to suppress the formation of the LSV.
To show the influence of the hypoviscosity on the flow we consider the kinetic energy spectrum
@@
Figure 1: (a) Instantaneous kinetic energy spectra $E(k)$ for the cases with $\textrm{Ra}=8\times 10^{6}$ with and without hypoviscosity and for different horizontal domain sizes (see table 1). (b,c) The corresponding snapshots of kinetic energy $\tfrac{1}{2}|\bm{u}|^{2}$ for the cases with $L/H=2.235$ at the mid-plane $z=H/2$.
@@
    The hypoviscosity needs to be able to dissipate the full inverse energy flux before it condenses into an LSV. In practice, this means that the hypoviscosity must be strong enough such that the energy contained in the smallest (system scale) wavenumber does not dominate the spectrum. This can be translated into a condition on the spectral slope at the smallest wavenumber, demanding that $\frac{\mathrm{d}\log E(k)}{\mathrm{d}\log k}\big|_{k=2\pi/L}>0$.
@@
    The action of the hypoviscosity needs to remain restricted to the smallest wavenumbers. We ensure this by checking that the spectral slope at small wavenumbers does not exceed $\frac{\mathrm{d}\log E(k)}{\mathrm{d}\log k}\big|_{k=2\pi/L}<2\alpha$, such that the action of the hypoviscosity is indeed dominated by the smallest wavenumbers.
Then, if there is sufficient scale separation between the large scales where the hypoviscosity acts and the scales that dominate the energy injection, conservation of energy ensures that the rate of energy dissipation by the hypoviscosity will match exactly with the inverse energy flux, independent of the precise values of the hypoviscosity parameters, as long as the aforementioned conditions are met. This is verified numerically for an exemplary case in appendix 7.
Under these conditions, fluctuation fields that are dominated by intermediate- and small-scale structures remain unaffected by the presence of hypoviscosity, see for example figure 2.
Figure 2: Instantaneous snapshots of the temperature fluctuation field $T-\tilde{T}(z)$, where $\tilde{T}(z)$ is the horizontally and temporally averaged temperature profile, for the cases with $\textrm{Ra}=8\times 10^{6}$ and $L/H=2.235$ without hypoviscosity (a) and with hypoviscosity (b).
@@ 4 Transition towards inverse cascade
With the hypoviscosity in place and tuned, we are well-equipped to study the transition towards the inverse energy cascade under the influence of large-scale friction, in absence of the LSV. To that extent, we consider the strength of the inverse energy flux $\varepsilon_{\textrm{inv}}$ as we vary the control parameter Ra throughout the regime where the inverse cascade is known to occur.
As argued above, when the action of the hypoviscosity is correctly tuned, the energy dissipated by the hypoviscosity coincides with the inverse energy flux. We can thus measure the mean inverse energy flux in this case as
$\varepsilon_{\textrm{inv}}^{\textrm{(hypo)}}\equiv\left\langle\nu_{\alpha}\bm{u}_{h}\cdot\Delta_{h}^{-\alpha}\bm{u}_{h}\right\rangle,$ (6)
where we use $\langle$$\rangle$ to denote the average over the full spatial domain and time.
@@
$\varepsilon_{\textrm{inv}}^{\textrm{(no-hypo)}}\equiv\sum_{Q}T_{\textrm{3D}}(K=\tfrac{2\pi}{L},Q)+T_{\textrm{2D}}(K=\tfrac{2\pi}{L},Q)$ (7)
@@
[…]
@@
Figure 3: The total inverse flux of kinetic energy $\varepsilon_{\textrm{inv}}$ normalized by the total energy flux $\varepsilon=(\textrm{Nu}-1)/\sqrt{\textrm{Pr}\textrm{Ra}}$ as a function of Ra for $\textrm{Ek}=10^{-4}$ and $\textrm{Pr}=1$ for the different series of runs with and without hypoviscosity. Panel (b) is a zoom of panel (a) that focuses on the cases with hypoviscosity, indicated by the dashed box in (a). For the runs without hypoviscosity, black diamonds indicate cases that show LSV formation. The figure reveals that while the transition to the inverse cascade without hypoviscosity is discontinuous and hysteretic (lower hysteretic branch is depicted by the dotted line), the transition for the case with hypoviscosity is continuous.
@@
The results are shown in figure 3. While the transition to the inverse cascade regime without large-scale friction, in presence of the LSV, is known to be discontinuous and hysteretic (Favier et al., 2019; de Wit et al., 2022a), we find that this behaviour is entirely different for the case with large-scale friction, in absence of the LSV. Here, we observe that the transition in this case is in fact smooth and the discontinuity disappears (and thereby also the hysteretic behaviour, see appendix 8). This is in agreement with what is found in other hypoviscous quasi-2D systems (Seshasayanan et al., 2014; Benavides & Alexakis, 2017; Pestana & Hickel, 2019; van Kan & Alexakis, 2020). We furthermore find that the strength of the inverse cascade in absence of the LSV is much smaller than for the cases that have an LSV.
@@
A point of attention is raised by the observation that the inverse cascade does not seem to decay to zero for large Ra for the case with hypoviscosity. However, we argue that this is an effect of insufficient scale separation between the large scales at which the hypoviscosity acts and the scales that dominate the energy injection, such that the hypodissipation not only measures the inverse energy flux, but also the fraction of energy that is directly injected at these large scales. The energy injection is known to be dominated by features of typical size $\mathcal{L}_{c}$, but the self-organizing nature of the forcing injects energy at a broad range of scales around this dominant scale. The runs without hypoviscosity (red line) were performed at $L\approx 10\mathcal{L}_{c}$ and the runs with hypoviscosity I (blue line) employ the same aspect ratio in order to have a one-to-one comparison. However, to increase the scale separation, we have added the series of runs in the wider domains $L\approx 20\mathcal{L}_{c}$ (II, green line) and $L\approx 40\mathcal{L}_{c}$ (III, yellow line), which also allows for a less aggressive exponent of the hypoviscosity (see table 1). Indeed, we notice that the measured energy flux that is dissipated by the hypoviscosity remains mostly unchanged, but reaches a smaller value at large Ra for the wider domains. We thus expect that the inverse flux will decay to values close to zero in the limit of increasingly large domain size as the scales become more cleanly separated.
@@
Figure 4: Time-averaged kinetic energy transport maps from 3D (a,c) and 2D (b,d) modes $Q$ to 2D modes $K$, i.e. $T_{\mathrm{3D}}(K,Q)$ and $T_{\mathrm{2D}}(K,Q)$, respectively, for the cases with $\textrm{Ra}=8\times 10^{6}$ and $L/H=2.235$ without hypoviscosity (a,b) and with hypoviscosity (c,d). Panels (a-d) also show the respective sums over the donating scales $Q$, obtained as $\mathcal{T}_{\mathrm{3D}}(K)\equiv\sum_{Q}T_{\mathrm{3D}}(K,Q)$ and $\mathcal{T}_{\mathrm{2D}}(K)\equiv\sum_{Q}T_{\mathrm{2D}}(K,Q)$ (blue lines). Panel (e) depicts the total 2D energy flux $\Pi_{\textrm{2D}}(K)\equiv-\sum_{K^{\prime}<K}\mathcal{T}_{\textrm{2D}}(K)$, normalized by the total energy flux $\varepsilon$.
@@
To further understand the influence of the presence/absence of the LSV on the nature of the inverse cascade, we also consider the scale-to-scale kinetic energy transfer maps. By looking at $T_{\textrm{3D}}(K,Q)$ and $T_{\textrm{2D}}(K,Q)$, we can study how baroclinic and barotropic kinetic energy is transferred from donating wavenumber $Q$ to receiving wavenumber $K$. The results are provided in figure 4 for an exemplary case with and without hypoviscosity. It shows that in presence of the LSV (upper panels), the upscale energy transfer is highly non-local, with a direct transfer from all wavenumbers $Q$ into the largest wavenumber $K=2\pi/L$, as is also known from Rubio et al. (2014); Aguirre Guzmán et al. (2020); de Wit et al. (2022a). In absence of the LSV (lower panels), on the contrary, we find that the upscale energy transfer has a much more local signature, with the energy transfer concentrating more around the diagonal of the map, indicating transfer between more like-sized scales. Note furthermore in figure 4e, that the total 2D energy flux $\Pi_{\textrm{2D}}(K)$ towards low $K$ for the case with hypoviscosity is found to be $\approx 0.003\varepsilon$, comparable to the total inverse energy flux measured by the hypodissipation (see figure 3b). This indicates that the inverse cascade that develops in the 2D manifold is fully dominating the measured total inverse flux in the case with hypoviscosity, while for the case without hypoviscosity (with LSV), the direct (non-local) baroclinic interactions between the large-scale 2D mode and the 3D modes are much stronger contributors (see figure 4a).
Finally, we can also study what the influence of the LSV is on the total thermal throughput of the convective system, quantified through the Nusselt number Nu, given by